# Quality Time: Carbon-Aware Quality Adaptation for Energy-Intensive Services

### v2

@@ 8. Conclusion
As data centers and cloud computing infrastructure continue to expand, their role in emissions reduction will be crucial for meeting global climate targets. In this paper, we explored the potential of adjusting a cloud service’s QoR to improve both energy and carbon efficiency. Our findings demonstrate that, in many regions, carbon-aware QoR adaptations can reduce emissions by up to 10 %, beyond the savings from reduced energy demand. Moreover, we propose an approach to automatically adapt the QoR of best-effort users to stay within a predefined annual carbon budget, which is an important contribution for service providers facing increasing pressure from carbon reporting and pricing.
@@ Acknowledgements.
This research was supported by the German Ministry for Education and Research (BMBF) as BIFOLD—Berlin Institute for the Foundations of Learning and Data (ref. 01IS18025A and 01IS18037A).
Appendix A Proof of Theorem 3.1
Theorem 3.1 We assume $p_{m,q}^{\text{max}}>p_{m}^{\text{idle}}>0$ and $n\in[0,1]$. When minimizing (10) subject to (11) – (14) from the perspective of data center operators, the utilization-dependent power model,
$p^{i}_{m,q}=p_{m}^{\text{idle}}+\bigl{(}p_{m,q}^{\text{max}}-p_{m}^{\text{idle}}\bigl{)}\cdot\bigl{(}\text{util}^{i}_{q}\bigl{)}^{n},$
can be simplified to
$p^{i}_{m,q}=p_{m,q}^{\text{max}},$
as it always yields an equivalent optimal deployment $D^{i}$.
Lemma A.1.
Increasing the utilization of machines does not result in additional penalties, so there is no incentive to deploy more machines than strictly necessary. Formally, if $p_{m,q}^{\text{max}}>p_{m}^{\text{idle}}>0$ and $n\in[0,1]$, an utilization-dependent power model $f$ is strictly subadditive
(18) $f(x+y)<f(x)+f(y),$
where $x$ and $y$ represent utilization levels $\text{util}^{i}_{q}\in(0,1]$.
Proof of Lemma A.1.
We show that $f$ is strictly subadditive for the two cases $n=1$ and $n<1$.
Case 1: Linear power model ($n=1$):
(19) $f(x+y)$ $=p_{m}^{\text{idle}}+\left(p_{m,q}^{\text{max}}-p_{m}^{\text{idle}}\right)\cdot\left(x+y\right)$
$f(x+y)$ $=p_{m}^{\text{idle}}+\left(p_{m,q}^{\text{max}}-p_{m}^{\text{idle}}\right)x$
(20) $\quad+p_{m}^{\text{idle}}+\left(p_{m,q}^{\text{max}}-p_{m}^{\text{idle}}\right)y-p_{m}^{\text{idle}}$
(21) $f(x+y)$ $=f(x)+f(y)-p_{m}^{\text{idle}}$
(22) $f(x+y)$ $<f(x)+f(y)$
Case 2: Sub-linear power model ($n<1$): For a strictly concave function $f$ with $f(0)\geq 0$, $f$ is strictly subadditive on $(0,\infty)$.
(23) $f(x)$ $=p_{m}^{\text{idle}}+\left(p_{m,q}^{\text{max}}-p_{m}^{\text{idle}}\right)\cdot x^{n}$
(24) $f^{\prime}(x)$ $=n\left(p_{m,q}^{\text{max}}-p_{m}^{\text{idle}}\right)\cdot x^{n-1}$
(25) $f^{\prime\prime}(x)$ $=\underbrace{n(n-1)\vphantom{\Bigl{(}}}_{<1}\ \underbrace{\left(p_{m,q}^{\text{max}}-p_{m}^{\text{idle}}\right)}_{>1}\cdot\underbrace{x^{n-2}\vphantom{\Bigl{(}}}_{>1}<0$
(26) $f^{\prime\prime}(x)<0$ $\implies f\text{ is strictly concave}$
Additionally, $f(0)=p_{m,q}^{\text{max}}>0$. ∎
Proof of Theorem 3.1.
For each interval $i$, we define the minimal deployment $\check{D}^{i}$ as any feasible solution that uses the smallest possible number of machines:
[…]
  1. (1)
    Feasibility with fewer machines is impossible. If a solution $\widehat{D}^{i}$ has fewer machines than $\check{D}^{i}$ in at least one dimension, it would violate Sufficient resources (12).
  2. (2)
    Using extra machines is suboptimal. By Lemma A.1, under the utilization-dependent power model, there is no incentive to deploy extra machines. Likewise, any extra machine increases the objective under the simplified power model.
We conclude that, for both power models, any optimal solution must yield an equivalent deployment $\check{D}^{i}$ for each interval $i$. ∎
Appendix B Request Traces
@@
This section describes all request trace datasets used in the experiments. The traces are visualized in Figure 12; Table 3 presents summary statistics.
@@
Figure 12. First month of data and KDE over all data points in 2023 for the eight datasets used in the evaluation.
@@
Appendix C Carbon Intensity Forecasts
@@
Table 4. Carbon forecast MAPE over 96 hours as reported in CarbonCast (Maji et al., 2023)

### v3

@@ 5. Conclusion
In this paper, we presented a first analysis of carbon-aware QoR adaptation, a strategy to reduce operational carbon emissions of cloud services without relying on geo-distributed load balancing or delay-tolerant workloads. Our results demonstrate that adjusting service quality in response to carbon intensity can reduce emissions by up to 10 % beyond savings related to increased energy efficiency.
To cover a wider range of realistic settings, future work should extend our analysis to scenarios with heterogeneous machines, multiple quality tiers, and diverse user groups. It is also important to examine user-side effects, such as whether low QoR might trigger behavioral responses like repeated requests, potentially offsetting carbon savings. Additionally, future work should explore the potential of QoR adaptation under sustainability metrics used for short-term demand management, like marginal carbon intensity or excess power during curtailments (Wiesner and Kao, 2025). These metrics exhibit higher temporal variability than average carbon intensity, potentially enabling higher savings and the use of shorter validity periods.
@@ Acknowledgements.
This research was supported by the German Ministry for Education and Research (BMBF) as BIFOLD—Berlin Institute for the Foundations of Learning and Data (ref. BIFOLD24B), the German Research Foundation (DFG) as FONDA—Foundations of Workflows for Large-Scale Scientific Data Analysis (ref. CRC 1404), and the EU Horizon project EXIGENCE under Grant Agreement No. 101139120.
Appendix A Nomenclature
Input variables
$\mathcal{U}$ Set of user groups $u$
$\mathcal{M}$ Set of machine types $m$
$\mathcal{Q}$ Tuple of two service quality tiers $q$
$\gamma\in\mathbb{N}$ Validity period length
$R^{i}\in\mathbb{R}^{\left|\mathcal{U}\right|}$ Number of requests $\left(r^{i}_{u}\right)$ for user group $u$ during  $i$.
$C^{i}\in\mathbb{R}$ Carbon intensity during $i$.
$P^{i}\in\mathbb{R}^{\left|\mathcal{Q}\right|\times\left|\mathcal{M}\right|}$ Power usage $\bigl{(}p^{i}_{m,q}\bigl{)}$ of machine $m$ serving quality $q$ during $i$.
$K\in\mathbb{R}^{\left|\mathcal{Q}\right|\times\left|\mathcal{M}\right|}$ Requests per interval $\bigl{(}k_{m,q}\bigl{)}$ that machine $m$ can serve at quality $q$.
$C_{m}^{\text{emb}}\in\mathbb{R}$ Attributed embodied emissions for running machine $m$ for time $\Delta$.
Decision variables
$D^{i}\in\mathbb{N}^{\left|\mathcal{Q}\right|\times\left|\mathcal{M}\right|}$ Deployment $\bigl{(}d^{i}_{m,q}\bigl{)}$, i.e., the number of machines $m$ running quality $q$ during $i$.
$A^{i}\in\mathbb{R}^{\left|\mathcal{U}\right|\times\left|\mathcal{Q}\right|}$ Number of requests $\bigl{(}a^{i}_{u,q}\bigl{)}$ served by quality $q$ for user group $u$ during $i$.
Appendix B Proof of NP-Hardness
We prove that the optimization problem defined in Equations (3)–(6) is NP-hard via a polynomial-time reduction from the Bin Packing problem, which is known to be NP-complete (Garey and Johnson, 1979). Bin Packing is defined as: Given a set of items with sizes $a_{j}\in(0,1]$ and an integer $B$, can the items be packed into at most $B$ bins of capacity 1?
Consider a simplified instance of our problem with a single interval $i$, one user group $u$, one machine type $m$, and one quality tier $q$. Each machine has capacity $k_{m,q}=1$. Requests correspond exactly to Bin Packing items, each with size $a_{j}$, and the total request volume is $\sum_{j}a_{j}$. In this setting, determining feasibility under Constraints (4) and (5) is equivalent to solving the Bin Packing problem: each machine corresponds to a bin, each request to an item, and the capacity constraint ensures that the sum of assigned request sizes per machine does not exceed 1. Since this special case of our problem is equivalent to Bin Packing, and Bin Packing is NP-complete, it follows that our optimization problem is NP-hard. $\square$
Appendix C Performance Analysis
To illustrate the performance improvements achieved by the proposed multi-horizon optimization, we analyze 240 experiments (all regions, all trace datasets, and $\text{QoR}_{\text{target}}\in\{0.3,0.5,0.7\}$) for the annual-horizon and daily-horizon optimization. These correspond to the long-term and short-term optimization components of our evaluated online approach.
Figure 5. CDF of optimization gap over time.
Figure 5 shows a CDF over the fraction of runs below a certain optimization gap over time. In the annual-horizon optimization, around 10 % of all runs have not found a solution within 1 % of the optimum within 10 minutes, while no run found an optimal solution within one hour. In the daily-horizon optimization, the median runtime was only 1.2 seconds and more than 80 % of all runs found an optimal solution within 10 seconds.
Appendix D Request Traces
@@
This section describes all request trace datasets used in the experiments. Table 3 presents summary statistics.
@@
Appendix E Carbon Intensity Forecasts
@@
Table 4. Carbon intensit forecast MAPE over 96 hours as reported in CarbonCast (Maji et al., 2023)
@@
Appendix F Impact of QoRtarget
Figure 6 shows the potential for relative savings across different $\text{QoR}_{\text{target}}$ and validity periods, for Wiki (en) in Germany and California. At $\text{QoR}_{\text{target}}=0$ and $\text{QoR}_{\text{target}}=1$, we serve all requests using either LLaMA 3.1 8B or 70B, leaving no flexibility for additional savings. A $\text{QoR}_{\text{target}}$ around 0.5 offers the greatest flexibility for carbon-aware QoR adaptation, resulting in the highest potential for carbon savings.
Appendix G Periods of Low Service Quality
While increasing $\gamma$ enhances potential for carbon savings, it also increases the probability of prolonged periods with low QoR, illustrated in Figure 7. For example, when optimizing for $\text{QoR}_{\text{target}}=0.5$ at $\gamma=1$ week, no 1-week interval has a QoR below 0.5 (right). However, 10 % of all daily intervals exhibit a $\text{QoR}_{\text{target}}<0.48$ (left). We observe that for $\gamma=3$ months, users can experience multiple consecutive days of low QoR. Consequently, large validity periods are not practical in many real-world scenarios.